OpenStudy (anonymous):

help please Consider the equation 7x + 3y = 42. Part 1: On your own paper, graph this equation using the slope-intercept method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar. Part 2: On your own paper, graph this equation using the intercepts method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar.

OpenStudy (anonymous):

part 1!! slope intercept form issssssssssss............ y = mx + b m is the slope, and b is the y intercept

OpenStudy (anonymous):

can you try put it in that form?

OpenStudy (anonymous):

7x+3y=42 3y=42-7x Or y=42-7x/3 y=14-7/3x So slope=-7/3 Also, y-intercept=14 Now we can plot the graph.  From the y-intercept 14, go down 7 units since it is -7/3 and 3 units towards the right in positive x-axis. You should get something like that.

OpenStudy (anonymous):

Now for the intercepts methods. Consider the equation y=-7/3x+14 Y-intercept is 14 right? Now find the x-intercept, to do so take F(x) or F(y)=0 so, 0=-7/3x+14 Take 14 to other side, -14=-7/3x Taking 3 to other side, -42=-7x - gets cancelled from both sides, so 42=7x 42/7=6=x So x-intercept=6